OpenStudy (anonymous):

What is the vertex of the parabola given by y = x^2 − 10x + 17?

OpenStudy (anonymous):

Im thinking it's (5,-8) ?

OpenStudy (anonymous):

Yup, good job! If you're given an equation in the form ax^2 +bx +c (remember that a, b, and c are just constants, they are just any integer), you can use the formula x= -b/2a to find the x-value of the vertex. In your case, b=-10, and a= 1. so if you plug it in, you'll get x=-(-10)/2(1). x=10/2. x=5. So 5 is the x-value of the vertex. To find the y-value (because your looking for a point), simply plug the x-value of 5 into the original equation y = x^2 − 10x + 17. you'll find that y=-8. Hope this helps and keep asking questions! :)

OpenStudy (anonymous):

Sure did! Thanks again :]

OpenStudy (anonymous):

anytime!